# Make Your AUV Adaptive: An Environment-Aware Reinforcement Learning Framework For Underwater Tasks

### v1

@@ III-A The Environment-Aware Module
In the proposed RL framework, the environment-aware module is responsible for introducing underwater flow field information into the RL training process, thereby enabling the AUV to perceive its surrounding environment during task execution. To better integrate this module into the RL algorithm, similar to the policy network and value network, we construct this module using a PINN [26] and name it the environment-aware network $\mathcal{U}_{E}(\mathbf{\vec{x}},t;\theta)$, where $\theta$ is the network’s parameter. Specifically, the network can output the corresponding flow field velocity $\mathbf{\vec{v}}$ and pressure $p$ based on the input position $\mathbf{\vec{x}}$ and time $t$.
@@
where $St$, $Eu$ and $Re$ represent the Strouhal number, Euler number and Reynolds number, respectively. $\mathcal{B}(\mathbf{\vec{x}}_{\mathcal{B}},t)$ represents the boundary conditions at boundary $\mathbf{\vec{x}}_{\mathcal{B}}$, and $\mathcal{I}(\mathbf{\vec{x}},t_{\mathcal{I}})$ denotes the initial conditions at the initial time $t_{\mathcal{I}}$.
@@ III-B The RL Training Module
=$\begin{split}\mathcal{M}=&\{[{\mathcal{S}}_{1}^{t},{\mathcal{S}}_{2}^{t},\ldots,{\mathcal{S}}_{N}^{t}],[{\mathcal{A}}_{1}^{t},{\mathcal{A}}_{2}^{t},\ldots,{\mathcal{A}}_{N}^{t}],\\ &[{\mathcal{R}}_{1},{\mathcal{R}}_{2},\ldots,{\mathcal{R}}_{N}],{\mathcal{P}}({s}^{t+1}|{s}^{t},{{a}}^{t}),\gamma\},\end{split}$ (8)
@@
State Space $\boldsymbol{\mathcal{S}}$: The state space in the environment-aware RL comprises two components: the AUV’s observed information $s^{{}^{\prime}}$, and features $f$ extracted from the flow field, represented as $\mathcal{S}=(s^{{}^{\prime}},f)\in\boldsymbol{S^{{}^{\prime}}}\times\boldsymbol{\mathcal{F}}$, where $\boldsymbol{S^{{}^{\prime}}}$ and $\boldsymbol{\mathcal{F}}$ denote the state intervals observed by the AUV and the flow field information interval, respectively. By incorporating flow field data into the state space, the AUV’s policy can adapt to varying flow conditions during training, facilitated by the environment-aware network based on PINN, as discussed in Section III-A. Additionally, we denote the AUV’s position in the coordinate system as $\boldsymbol{\vartheta}^{t}=[p_{x}^{t},p_{y}^{t},p_{z}^{t}]^{T}$ to represent the position of the AUV in the coordinate system, where $(p_{x}^{t},p_{y}^{t},p_{z}^{t})$ represent the AUV’s horizontal, longitudinal, and depth positions, respectively.
@@
Reward Function $\boldsymbol{\mathcal{R}}$: In the MDP, the AUV evaluates action policies based on environmental rewards. During environment-aware RL training, as the AUV adapts to the underwater flow field, task completion improves compared to traditional RL frameworks, leading to higher rewards. The reward function typically includes components like failure penalties, efficiency promotion, and cooperation encouragement, with its design varying depending on the specific task.
@@ III-C LLM-Based AUV Structure Optimization Module
[…]
@@
Length of AUVs 5 $\mathrm{m}$
Width of AUVs 1.5 $\mathrm{m}$
@@
Training episodes 300
The learning rate 1×$\times 10^{-4}$
@@ IV-A Experiment Setup
Multi-AUV data collection task: This task involves the coordinated operation of two AUVs in a 200 m×$\times$200 m×$\times$200 m area. to maximize data collection rates by prioritizing data from urgent nodes while minimizing energy consumption and avoiding collisions. Two experimental scenarios are established: normal and complex sea conditions, to assess the framework’s robustness against environmental disturbances. For further details on this task, due to the constraint of space, please refer to this previous work [28].
@@ IV-B Experiment I: Multi-AUV Data Collection Task Under Normal Sea Conditions
We first applied environment-aware RL framework to train two AUVs to complete the data collection task under normal sea conditions, excluding complex disturbances like tides and turbulence. The AUVs utilized the embedded environment-aware module to capture near-field hydrodynamic data, which was integrated into the state space representation. After each training session, the LLM-based AUV structure optimization module quantitatively assessed the performance metrics in the training results and dynamically adjusted AUV morphology. Our experiment involved three iterative design generations, each consisting of 300 RL training episodes, whose results are shown in Fig. 4.
Fig. 4 presents a comparative analysis of training results of environment-aware RL versus the traditional RL framework. Results show that the environment-aware RL framework consistently outperforms the traditional RL framework across all AUV generations. Both cumulative reward and data rate improved continuously during training, indicating the development of adaptive control policy for environmental compliance and task objectives. Additionally, energy consumption stabilized quickly despite performance gains, demonstrating the proposed framework’s energy efficiency. Evolving structure designs yielded significant improvements in reward and data rate while reducing energy consumption, suggesting that optimized AUV’s structure enhance environmental adaptability through reduced drag and improved flow field. The details of AUV structure and the structure of each AUV generation with corresponding flow field information are presented in Table II and Fig. 5, respectively.
Fig. 5 depict the design refinements across three AUV generations, along with their corresponding flow field environments. The initial capsule-structured design offers stability in underwater navigation but is suboptimal for complex tasks. Using RL training results and flow field data from the 1st generation (1-gen), the 2nd generation (2-gen) AUV was redesigned with an ”ice cream cone” structure, improving flow distribution and reducing drag while enhancing maneuverability and stability. While the 2nd generation features a linear conical structure, the 3rd generation (3-gen) adopts a parabolic tail for further drag reduction. Additionally, the transition between the semicircular head and conical body in the 3rd generation is smoother, minimizing turbulence and enhancing the AUV’s efficiency and task completion capabilities. Table II compares the structural designs of three generations of AUVs, where $\zeta=\frac{x-(C_{x}-L/2)}{L}$, $C_{x}$ and $C_{y}$ denote the horizontal and vertical coordinate of the central point, and $W$ and $L$ represent the width and length of the AUV, respectively.

### v2

@@ III-A The Environment-Aware Module
In the proposed RL framework, the environment-aware module is responsible for introducing underwater flow field information into the RL training process, thereby enabling the AUV to perceive its surrounding environment during task execution. To better integrate this module into the RL algorithm, similar to the policy network and value network, we construct this module using a PINN [RAISSI2019686] and name it the environment-aware network $\mathcal{U}_{E}(\mathbf{\vec{x}},t;\theta)$, where $\theta$ is the network’s parameter. Specifically, the network can output the corresponding flow field velocity $\mathbf{\vec{v}}$ and pressure $p$ based on the input position $\mathbf{\vec{x}}$ and time $t$.
@@
where $St$, EuEu and $Re$ represent the Strouhal number, Euler number and Reynolds number, respectively. $\mathcal{B}(\mathbf{\vec{x}}_{\mathcal{B}},t)$ represents the boundary conditions at boundary $\mathbf{\vec{x}}_{\mathcal{B}}$, and $\mathcal{I}(\mathbf{\vec{x}},t_{\mathcal{I}})$ denotes the initial conditions at the initial time $t_{\mathcal{I}}$.
@@ III-B The RL Training Module
$\begin{split}\mathcal{M}=&\{[{\mathcal{S}}_{1}^{t},{\mathcal{S}}_{2}^{t},\ldots,{\mathcal{S}}_{N}^{t}],[{\mathcal{A}}_{1}^{t},{\mathcal{A}}_{2}^{t},\ldots,{\mathcal{A}}_{N}^{t}],\\ &[{\mathcal{R}}_{1},{\mathcal{R}}_{2},\ldots,{\mathcal{R}}_{N}],{\mathcal{P}}({s}^{t+1}|{s}^{t},{{a}}^{t}),\gamma\},\end{split}$ (8)
@@
[…]
@@
[…]
@@ III-C LLM-Based AUV Structure Optimization Module
$\pi^{*}=\underset{{f}}{\mathrm{argmax}}\underset{T\rightarrow\infty}{\mathrm{lim}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{{R}}\left(\pi,f\right)\right].$ (9)
@@
Length of AUVs 1.5 $\mathrm{m}$
Width of AUVs 0.5 $\mathrm{m}$
@@
Training episodes 400
[…]
@@ IV-A Experiment Setup
Multi-AUV data collection task: This task involves the coordinated operation of two AUVs in a $\times200 m$$\times$200 m area. to maximize data collection rates by prioritizing data from urgent nodes while minimizing energy consumption and avoiding collisions. Two experimental scenarios are established: normal and complex sea conditions, to assess the framework’s robustness against environmental disturbances. For further details on this task, due to the constraint of space, please refer to this previous work [38].
@@ IV-B Experiment I: Multi-AUV Data Collection Task Under Normal Sea Conditions
We first applied environment-aware RL framework to train two AUVs to complete the data collection task under normal sea conditions, excluding complex disturbances like tides and turbulence. The AUVs utilized the embedded environment-aware module to capture near-field hydrodynamic data, which was integrated into the state space representation. After each training session, the LLM-based AUV structure optimization module quantitatively assessed the performance metrics in the training results and dynamically adjusted AUV morphology. Our experiment involved three iterative design generations, each consisting of 400 RL training episodes (results from the first 300 training episodes), whose results are shown in Fig. 4.
Fig. 4 presents a comparative analysis of training results of environment-aware RL versus the traditional RL framework. Results show that the environment-aware RL framework consistently outperforms the traditional RL framework across all AUV generations (notably in early performance gains and convergence speed). Both cumulative reward and data rate improved continuously during training, indicating the development of adaptive control policy for environmental compliance and task objectives. Additionally, energy consumption stabilized quickly despite performance gains, demonstrating the proposed framework’s energy efficiency. Evolving structure designs yielded significant improvements in reward and data rate while reducing energy consumption, suggesting that optimized AUV’s structure enhance environmental adaptability through reduced drag and improved flow field. The details of AUV structure and the structure of each AUV generation with corresponding flow field information are presented in Table II and Fig. 5, respectively.
Fig. 5 depict the design refinements across three AUV generations, along with their corresponding flow field environments. As a schematic illustration, Fig. 5 presents flow fields around precise geometric configurations of three generations of AUV, simulated via transient Navier-Stokes equations using ANSYS Fluent. The initial capsule-structured design offers stability in underwater navigation but is suboptimal for complex tasks. Using RL training results and flow field data from the 1st generation (1-gen), the 2nd generation (2-gen) AUV was redesigned with an ”ice cream cone” structure, improving flow distribution and reducing drag while enhancing maneuverability and stability. While the 2nd generation features a linear conical structure, the 3rd generation (3-gen) adopts a parabolic tail for further drag reduction. Additionally, the transition between the semicircular head and conical body in the 3rd generation is smoother, minimizing turbulence and enhancing the AUV’s efficiency and task completion capabilities. Table II compares the structural designs of three generations of AUVs, where $\zeta=\frac{x-(C_{x}-L/2)}{L}$, $C_{x}$ and $C_{y}$ denote the horizontal and vertical coordinate of the central point, and $W$ and $L$ represent the width and length of the AUV, respectively.
@@ V CONCLUSION
VI ACKNOWLEDGEMENT
The authors gratefully acknowledge the anonymous reviewers for their constructive comments. We also extend our sincere thanks to Dr. Xiangwang Hou, Prof. Yong Ren, Prof. Daoyi Chen, and Prof. Juntian Qu from Tsinghua University for their insightful discussions and guidance. Additionally, we appreciate the encouragement and recognition from Prof. Nare Karapetyan at the Woods Hole Oceanographic Institution, Prof. Xiaofan Li at the University of Hong Kong, and Prof. Xiaomin Lin at the University of South Florida.